OpenStudy (anonymous):

Select the equations that are parallel and perpendicular to y = -3x - 1 and that pass through the point (3, 1). parallel: y = -3x + 6 perpendicular: y = 1/3x + 2 2/3 parallel: y = -3x + 10 perpendicular: y = 1/3x parallel: y = negative 1/3x perpendicular: y = -3x parallel: y = 1/3x + 1 perpendicular: y = 3x - 1

OpenStudy (anonymous):

well parallel means same slope perpendicular means product of slopes=-1

OpenStudy (aravindg):

two lines having slopes m1 and m2 are parallel if m1=m2 perpendicular if m1 X m2 =-1

OpenStudy (anonymous):

I still do not get it.

OpenStudy (anonymous):

When two lines are parallel, they have the same slope. Example: \[y=-\frac{1}{2}x+4\text{ and }y=-\frac{1}{2}x-9\] are parallel, because their slopes are both 1/2. When two lines are perpendicular, their slopes are negative reciprocals of each other. Example: \[y=-\frac{1}{2}x+4\text{ and }y=2x-9\] are perpendicular, since the negative reciprocal of -1/2 is -(-2/1) = 2. Along with parallel/perpendicular, you have to choose a line that passes through (3,1). First, find the slope of the given line. Then, use the following point-slope formula: \[y-y_0=m(x-x_0),\text{ where $(x_0,y_0)$ is the given point and $m$ is the slope.} \]

OpenStudy (anonymous):

Which answer would it be?